Using Snell’s law, determine the constant when the angle of incidence and the angle of refraction are:


a) 50˚ and 30˚


b) 30˚ and 18˚


c) 60˚ and 38˚

1
Expert's answer
2022-05-31T10:05:16-0400

"n=\frac{\sin\alpha}{\sin\beta},"

a)

"n_1=1.53,"

b)

"n_2=1.62,"

c)

"n_3=1.41."